Calculate the net income earned during the year. Assume that the change to stockholder's equity results only from net income earned during the year.

Balance Jan. 1, 2019
Asset: $50,000
Liabilities: $40,000

Balance Dec 31, 2019
Assets: $35,000
Liabilities: $20,000
Answer #1

Total assets=Total liabilities+Total equity

Total equity at beginning=(50,000-40,000)=$10,000

Total equity at end=(35000-20,000)=$15,000

Ending equity=Beginning equity+Net income

Net income=(15000-10,000)

=$5,000